OpenStudy (anonymous):

Graph -1/4t is more than or equal to 9. Does the solution set get ( or [

OpenStudy (texaschic101):

-1/4t >= 9 -- multiply by 4 -t >= 36 t <= - 36 or -1/4t >= 9 (divide both sides by -1/4) t <= 9/(-1/4) t <= 9 * -4/1 t <= - 36 no ( or [ is needed

OpenStudy (anonymous):

\[\frac{-1}{4}t \ge9\]\[\frac{1}{4}t \le-9\]\[t \le-36\]\[(-\infty ,-36]\]